OpenStudy (anonymous):

graph each equation by writing two linear equations. y= |4x+3|

OpenStudy (ranga):

First we need to find when 4x+3 equals 0. 4x+3 = 0 4x = -3 x = -3/4. So when x < -3/4, 4x + 3 will be negative. Therefore, y = -(4x + 3) = -4x - 3 When x > -3/4, 4x + 3 will be positive. Therefore, y = +(4x + 3) = 4x + 3 Those are the two linear equations: y = -4x - 3 for x < -3/4 y = 4x + 3 for x > -3/4

OpenStudy (ranga):